An analyst has developed the following probability distribution for the rate of return for a common stock. f Return 0.29 0.50 0.21 -17% 158 a. Calculate the expected rate of return. (Round intermediate calculations to at least 4 decimal places. Round your answer to 2 decimal places.) Expected rate of return b. Calculate the variance and the standard deviation of this probability distribution. (Use the percentage values for your calculations (for example 10% not 0.10). Round intermediate calculations to at least 4 decimal places. Enter your answer as a percentage rounded to 2 decimal places) Variance Standard deviation
Answer #1

(a)

Expected rate of return, E(X) = 0.29 * (-17) + 0.5 * 7 + 0.21 * 15 = 1.72 %

(b)

Variance = E(X2) - [E(X)]2 = 0.29 * 172 + 0.5 * 72 + 0.21 * 152 - 1.722 = 152.6016

Standard Deviation = V152.6016 12.3532

** If the answer does not match please comment